LÉVY PHOTON STATISTICS IN BLINKING QUANTUM DOTS

\underline{ELI BARKAI}, Dept. of Chemistry, Notre Dame University, Notre Dame, IN 46556; e-mail: jbarkai@nd.edu; YOUNJOON JUNG, Dept. of Chemistry, University of California, Berkley, CA 94720-1460; BOB SILBEY, Dept. of Chemistry, MIT, 77 Mass. Ave., Cambridge, MA 02139.

We discuss the theory of blinking quantum dots.^{a} These systems exhibit non–ergodic, and non-stationary kinetics. The mean on and off mean off time diverge. We show that Mandel’s Q parameter, describing the fluctuations in this system, exhibits a new type of behavior, it increases with measurement time, even in the long time limit. In this case photon counts are described using Lévy statistics. Possible explanations of the power law kinetics are discussed.

aY. Jung, E. Barkai, and R. Silbey, Lineshape Theory and Photon Counting Statistics for Blinking Quantum Dots: a Lévy Walk Process Chemical Physics 284 181 (2002).